# HoneyGPT: Breaking the Trilemma in Terminal Honeypots with Large Language Model

### v1

@@
Abstract.
Honeypots, as a strategic cyber-deception mechanism designed to emulate authentic interactions and bait unauthorized entities, continue to struggle with balancing flexibility, interaction depth, and deceptive capability despite their evolution over decades. Often they also lack the capability of proactively adapting to an attacker’s evolving tactics, which restricts the depth of engagement and subsequent information gathering. Under this context, the emergent capabilities of large language models, in tandem with pioneering prompt-based engineering techniques, offer a transformative shift in the design and deployment of honeypot technologies. In this paper, we introduce HoneyGPT, a pioneering honeypot architecture based on ChatGPT, heralding a new era of intelligent honeypot solutions characterized by their cost-effectiveness, high adaptability, and enhanced interactivity, coupled with a predisposition for proactive attacker engagement. Furthermore, we present a structured prompt engineering framework that augments long-term interaction memory and robust security analytics. This framework, integrating thought of chain tactics attuned to honeypot contexts, enhances interactivity and deception, deepens security analytics, and ensures sustained engagement.
The evaluation of HoneyGPT includes two parts: a baseline comparison based on a collected dataset and a field evaluation in real scenarios for four weeks. The baseline comparison demonstrates HoneyGPT’s remarkable ability to strike a balance among flexibility, interaction depth, and deceptive capability. The field evaluation further validates HoneyGPT’s efficacy, showing its marked superiority in enticing attackers into more profound interactive engagements and capturing a wider array of novel attack vectors in comparison to existing honeypot technologies.
1. Introduction
Honeypots have been widely used as effective cybersecurity tools for detecting, enticing, and understanding malicious activities on the Internet  (López-Morales et al., 2020; Kim et al., 2022; Sasaki et al., 2022; Anagnostakis et al., 2005; Wahab et al., 2019; Rrushi, 2018; Torabi et al., 2020; Zhan et al., 2013; Wang et al., 2023; Huang and Zhu, 2021). These traps are designed to lure attackers, allowing security teams to gather their information and monitor their behaviors for threat analysis, as well as
deflect ongoing attacks.
Terminal honeypots are a specialized type of honeypots for emulating real terminal systems. They are typically employed to attract hackers or malicious software targeting terminal systems. Terminal honeypots range from low-interaction variants that simulate a minimal set of protocols or service information  (Niels, 2004), to medium-interaction honeypots that replicate a broader array of system functionalities at a higher deployment and maintenance cost, and high-interaction honeypots that leverage sophisticated virtualization technologies, such as VMWare to create convincing environments of terminal systems or even deploy actual systems for maximum authenticity.
@@ 1. Introduction
The development of Large Language Models (LLMs) (Kaplan et al., 2020; Wei et al., 2022a; Fu et al., 2022; Feng et al., 2024; Yang et al., 2022b) appears to offer a potential solution to these problems. By providing attack commands as queries to the large language model, it can generate responses tailored to the attackers’ interests based on the provided prompt. This approach breaks the Trilemma of traditional honeypots, as a single LLM can simultaneously simulate multiple high-interaction honeypots with different configurations. Furthermore, by designing prompts to guide the LLM, it is possible to better fulfill attackers’ objectives and entice them into deeper levels of interaction.
However, designing honeypots based on LLMs still faces three major challenges. A prominent issue is the absence of a standardized approach to honeypot prompt design. To address this problem, this paper establishes a universal specification for honeypot prompt keywords. Another challenge arises from the native thinking constraints of LLMs, which cannot handle complex attack combinations. To address this challenge, we employ a Chain of Thought (CoT) strategy that allows the language model to answer the command’s impact on the operating system during each interaction. When posing the next question, we integrate the consolidated changes in the operating system and the user command to enhance the large language model’s understanding of the task. The third challenge lies in that the performance of LLMs in prolonged sessions is hampered by their context length limitations and inherent forgetting mechanisms. To manage this issue, we employ the CoT approach to assess the significance of each attacker command and its impact on the terminal system. This assessment allows us to strategically prune the session history, ensuring that only the most essential contextual information is retained within the confines of the context length, thereby optimizing the honeypot’s effectiveness in extended interactions. The major contributions of this paper are summarized as follows.
@@
2. We propose a Chain of Thought strategy for terminal honeypot task scenarios, which enhances the task-solving capabilities of native LLMs and overcomes the inherent context length limitations. The introduction of CoT is essential to advancing honeypot technology and improving network security defense capabilities.
@@
4. We deploy HoneyGPT and Cowire, which is a programmatic honeypot, on the Internet for four weeks and observe that HoneyGPT captures more attack behaviors and longer interactions than Cowrie.
The rest of this paper is structured as follows. Section 2 provides background on LLMs, prompt engineering, and terminal honeypots. Section 3 describes the trilemma of flexibility, interaction, and deception in existing terminal honeypots. Section 4 presents our HoneyGPT solution, detailing its framework, concept, strategy, prompt management, and configuration. Section 5 validates the efficacy of HoneyGPT, providing baseline comparisons on deception, interaction, and flexibility, as well as field evaluations to measure its effectiveness in engaging attackers and capturing novel attack vectors. Section 6 outlines the limitations and potential future work, and finally, Section 7 concludes the paper.
2. Background
@@
2.1. Large Language Models
Large Language Models (LLMs) are a subset of pre-trained language models (PLMs), characterized by their voluminous parameter sets, reaching or even exceeding hundreds of billions, as defined by Kaplan et al (Kaplan et al., 2020). Since OpenAI introduced ChatGPT  (openai, [n. d.]), LLMs have been widely used across a spectrum of disciplines  (Fu et al., 2022; Brown et al., 2020), particularly in the field of cybersecurity  (Meng et al., 2024; Hu et al., 2024; He and Vechev, 2023; Pearce et al., 2023, 2022; Vishwamitra et al., 2024). Compared to previous models, LLMs offer two main advantages in interdisciplinary applications:
Firstly, LLMs display astonishing emergent abilities as identified by Wei et al (Wei et al., 2022a), which are competencies absent in smaller-scale models but manifest in larger configurations. Such models experience a quantum leap in performance relative to their diminutive counterparts. Prominent among these emergent abilities are:
  • In-context learning: LLMs can, upon receiving natural language instructions or several task demonstrations, generate accurate outputs for test cases by continuing the input text sequence, circumventing the need for further training or gradient adjustments  (Brown et al., 2020).
  • Instruction following: LLMs possess the ability to comprehend and execute natural language instructions with minimal or even zero examples, adapting to new tasks (Victor et al., 2022). By fine-tuning tasks articulated through instructional prompts, LLMs have demonstrated proficiency in accurately handling tasks they have never seen before (Ouyang et al., 2022).
  • Step-by-step reasoning: Distinct from conventional PLMs, LLMs are capable of deconstructing complex tasks into a sequence of discrete subtasks by employing prompt-based strategies such as chain-of-thought. It is conjectured that this capacity may be derived from their training on code  (Wei et al., 2022b; Fu et al., 2022).
@@ 2.2. Prompt Engineering
Components
The standard components of LLM prompts typically include:
  • Task Description: Instructions that the LLM is expected to adhere to, expressed in natural language.
  • Input Data: The requisite data for a task, presented in natural language.
  • Contextual Information: The contextual or background information that can aid in the execution of a task.
  • Prompt Style: The formulation of the prompt, which is tailored to optimize the model’s responses, such as through role-playing or incremental reasoning.
Design Principles
The development of effective LLM prompts is predicated on:
@@
  • Clarity: Ensure instructions are explicit and communication is clear through symbolic delimiters, structured outputs, conditional logic, and ”Few-shot” examples.
  • Deliberation: Allow LLMs time for thought, structuring prompts for sequential processing and reasoning with ”chain of thought” techniques. Allow LLMs time to reason by structuring prompts for sequential information processing using ”chain of thought” techniques.
Prompting Strategies
@@
The ”Chain of Thought” (CoT)  (Wei et al., 2022b; Feng et al., 2024; Yang et al., 2022b; Ling et al., 2024; Lu et al., 2022; Yang et al., 2022a; Zelikman et al., 2022) paradigm has been shown to significantly enhance LLM performance on tasks that require complex reasoning by incorporating intermediary reasoning steps within the prompts. The zero-shot CoT, as introduced in foundational research, employs prompts such as ”Let’s think step by step” to shepherd LLMs to a logical conclusion, thus highlighting the significance of scale in emergent capabilities.
2.3. Terminal Honeypots
@@
Honeypots are security tools designed as decoys to attract and monitor unauthorized or malicious attackers. Based on their implementation approaches, terminal honeypots can be categorized into three primary types: emulated honeypots, which are based on programmed simulations; real system honeypots, which utilize real systems; and hybrid honeypots, which combine the features of the first two types.
Emulated Honeypots
Emulated honeypots are software constructs that simulate network services and devices to attract cyber adversaries. They are developed using conventional programming methodologies, as exemplified by Honeyd  (Niels, 2004), Dionaea (phibos, [n. d.]), and Cowrie (OOSTERHOF, [n. d.]). These honeypots utilize rule-based matching and logical controls to mimic desired targets. Despite their effectiveness, they are constrained by the limitations of development and operational costs, preventing a full emulation of an operating system’s intricate behaviors. The interactivity spectrum of these honeypots spans from minimal to moderate: Honeyd and Dionaea deliver low-interaction platforms rich in configuration possibilities, while Cowrie steps up to offer a medium-interaction milieu. However, this comes at the cost of diminished adaptability and increased operational costs. Such frameworks have laid the groundwork for a range of specialized offshoots (López-Morales et al., 2020; Alexander and Richard, 2018).
Real-system Honeypots
Real-system honeypots, by contrast, utilize genuine, unmodified system environments, offering a more authentic simulation of user interactions, file system activities, and operational behaviors  (Wang et al., 2022; Anagnostakis et al., 2005; Wahab et al., 2019). They provide a high level of interaction, creating a convincing facade of a real system, which significantly increases the probability of capturing complex attack strategies that simpler honeypot architectures may miss. However, the inherent rigidity of their physical environments (Alexander and Richard, 2018) requires individualized configuration for each new system variant, which incurs substantial deployment and maintenance costs.
This authenticity offers potential attackers a highly interactive environment, fostering a convincing deception that aids in uncovering intricate attack vectors that might elude detection by more rudimentary honeypots. However, a challenge arises from the security protocols native to real operating systems, such as stringent permission controls, which are not inherently conducive to fulfilling malicious requests. These protocols can inadvertently deter threat actors from engaging in more profound levels of system exploitation, thus potentially curtailing the honeypot’s effectiveness as a lure. Additionally, the deployment and maintenance of real-system honeypots are encumbered by the need for individual configurations for each system variant, leading to significant resource expenditures.
Hybrid Honeypots
In an effort to combine the strengths of both emulated and real system honeypots, a scalable hybrid honeypot framework has been proposed (Xu et al., 2021; You et al., 2021). This framework employs programmed methods for simulating low-level interactions, such as network protocols, while delegating high-level interactions, such as file executions, to a real file system. Although the hybrid model ostensibly enhances flexibility and interaction depth, it is not devoid of limitations. The framework may experience discordance between the simulated protocols and the actual file system responses, leading to a detectable discrepancy that astute attackers could potentially exploit. Furthermore, the rigidity in configuring a physical environment restricts adaptability and diminishes the diversity of supported environmental configurations, which could impair its capacity to effectively deceive sophisticated attackers.
3. Trilemma of Existing Terminal Honeypots
@@
The trilemma in honeypot design reflects the intricate balancing act required to optimize the three key attributes of effective honeypot systems: interaction depth, flexibility, and deception ability. Honeypots from their inception have grappled with this trilemma, consistently facing trade-offs that impede their overall efficacy.
3.1. Limited Flexibility
The narrow concept of flexibility within honeypot design is indicative of the system’s adaptability in authentically replicating a variety of systems and services. Such adaptability is pivotal for honeypots, as it underpins their ability to create authentic and persuasive simulations of various operational environments. Yet, the quest for such flexibility is one of the most formidable challenges in honeypot development, with the central issue being the support for a wide range of operating system versions and nuanced system configurations.
Contemporary honeypots often struggle to provide sufficient flexibility, particularly regarding the support of diverse operating system versions and security configurations. The pursuit of increased flexibility can compromise the depth of interaction and the honeypot’s deception ability.
Code-based emulated honeypots exemplify this issue, as they require extensive resources for the creation and maintenance of simulations that encompass numerous operating system versions and environmental configurations. Such complexity leads to substantial development and operational costs, and no code-based honeypot has yet managed to seamlessly transition across varied terminal environments while maintaining high interaction fidelity. Real-system honeypots offer detailed interactions but lack the adaptability of their emulated counterparts, making them difficult to reconfigure in response to evolving threats. They also impose significant management overhead, due to the requirement of maintaining real operating system environments.
Moreover, traditional honeypot offerings remain constrained to single-dimensional deception techniques. As to leveraging the data they capture, there is a conspicuous lack of flexibility in integrating security analysis, necessitating distinct systems dedicated to data processing and analysis.
The core challenge, therefore, is to architect honeypots that strike an optimal balance between delivering authentic, complex interactions and possessing the agility to counter novel attacks. Achieving this equilibrium should not exacerbate costs or inflate management burdens, crafting a solution that is as economical as it is effective.
3.2. Limited interaction
Due to limitations in implementation methods and development costs, code-based terminal honeypots are restricted to simple rule matching and lack the capability to execute complex commands as a real operating system. This limitation significantly reduces the level of interaction these honeypots can manage, making them inadequate in responding to complex commands issued by attackers. For instance, Cowrie, a commonly used middle-interactive honeypot, is unable to handle complex commands that involve concatenation with pipe symbols (—). The inability to process complex operations can lead to honeypot instability or crashes, thereby failing to further engage attackers. Enhancing the interaction of honeypots is crucial for increasing their reliability and effectiveness.
3.3. Limited Deception
A key factor contributing to limited deception is the lack of dynamic and adaptive behavior in existing honeypot systems. Code-based terminal honeypots typically rely on static configurations and predefined matching mechanisms, while honeypots based on real operating systems often perform limited and monotonous simulations of the operating environment due to cost constraints. Experienced attackers can easily identify these singular and rigid environment configurations. The lack of variability reduces the effectiveness of deception because attackers can quickly recognize and bypass the simulated environment of the honeypot. Additionally, code-based terminal honeypots lack authentic user interactions and the realism of contextual information, further limiting their ability to convincingly simulate real systems. When attackers perform complex operations such as file execution or process communication, code-based honeypots struggle to replicate authentic user interactions or provide the depth of a real system.
In addition, in terms of enticing attackers for deeper-level attacks, current honeypots often fall short. Whether it is a code-based honeypot like Cowrie or a honeypot based on real systems, the limited configurations and restricted permissions provided to attackers often fail to meet the intentions of attackers. Catering actions such as modifying user passwords, terminating competing processes, or viewing graphics card information, which could effectively boost an attacker’s motivation, are often overlooked by existing honeypots. Especially in a real system environment, the operating system’s characteristics, such as permission control mechanisms and limited configurations (such as hardware devices and account password configurations), have limitations in fulfilling the attacker’s attack intentions.
4. HoneyGPT
@@
To address the trilemma of existing honeypots, we introduce HoneyGPT, an advanced honeypot framework based on ChatGPT. This framework transforms the conventional ”request-response” message interaction (based on terminal protocols) into a ”question-answer” text interaction (based on ChatGPT API).
@@
As shown in Figure  1, the framework consists of three main components: the Terminal Protocol Proxy, the Prompt Manager, and ChatGPT. The Terminal Protocol Proxy is responsible for fundamental protocol tasks, such as establishing connections, fingerprint emulation, message encapsulation, and parsing. We utilize the protocol layer of the Cowrie honeypot to develop the Terminal Protocol Proxy, which supports protocols such as SSH and Telnet. The Prompt Manager is tasked with constructing prompts based on commands sent by attackers and interacting with ChatGPT. ChatGPT generates responses based on the provided prompts. After receiving a response from ChatGPT, the Prompt Manager extracts the terminal output and forwards it to the Terminal Protocol Proxy, which then encapsulates it into messages and returns them to attackers. Since the Terminal Protocol Proxy repurposes the work of the Cowrie honeypot and is not a primary innovation, we will predominantly elaborate on the design of the Prompt Manager in the following section.
Figure 1. HoneyGPT Framework
Figure 2. HoneyGPT Architecture
4.1. Concept and Strategy
We first articulate the fundamental concepts and methodological strategies that are integral to the design of the Prompt Manager.
4.1.1. Concept of HoneyGPT

### v2

# HoneyGPT: Breaking the Trilemma in Honeypots with Large Language Models

@@
Abstract
Honeypots, as a strategic cyber-deception mechanism designed to emulate authentic interactions and bait unauthorized entities, often struggle with balancing flexibility, interaction depth, and deception. They typically fail to adapt to evolving attacker tactics, with limited engagement and information gathering. Fortunately, the emergent capabilities of large language models and innovative prompt-based engineering offer a transformative shift in honeypot technologies. This paper introduces HoneyGPT, a pioneering shell honeypot architecture based on ChatGPT, characterized by its cost-effectiveness and proactive engagement. In particular, we propose a structured prompt engineering framework that incorporates chain-of-thought tactics to improve long-term memory and robust security analytics, enhancing deception and engagement. Our evaluation of HoneyGPT comprises a baseline comparison based on a collected dataset and a three-month field evaluation. The baseline comparison demonstrates HoneyGPT’s remarkable ability to strike a balance among flexibility, interaction depth, and deceptive capability. The field evaluation further validates HoneyGPT’s superior performance in engaging attackers more deeply and capturing a wider array of novel attack vectors.
I Introduction
Honeypots have been widely used as an effective approach for detecting, enticing, and understanding malicious activities on the Internet  [14, 12, 25, 2, 31, 24, 28, 41, 32, 10].
These traps deceive attackers, enabling defenders to gather information, monitor behaviors for threat analysis, and prevent ongoing attacks. Terminal honeypots are a specialized type of honeypots for emulating real terminal systems to attract hackers or malicious software targeting terminal systems. Terminal honeypots range from low-interaction variants that simulate a minimal set of protocols or service information  [17], to medium-interaction honeypots that replicate a broader array of system functionalities at a higher deployment and maintenance cost, and high-interaction honeypots that leverage sophisticated virtualization technologies, such as VMWare to create convincing environments of terminal systems or even deploy actual systems for maximum authenticity.
@@ I Introduction
The development of Large Language Models (LLMs) [11, 34, 6, 5, 38] appears to offer a potential solution to these problems. By providing attack commands as queries to the large language model, it can generate responses tailored to the attackers’ interests based on the provided prompt. This approach breaks the trilemma of traditional honeypots, as a single LLM can simultaneously simulate multiple high-interaction honeypots with different configurations. Furthermore, by designing prompts to guide the LLM, it is possible to better fulfill attackers’ objectives and entice them into deeper levels of interaction.
However, designing honeypots based on LLMs still faces three major challenges. A prominent issue is the absence of a standardized approach to honeypot prompt design. To address this problem, this paper establishes a universal specification for honeypot prompt keywords that adapts to long dialogues and enables real-time updates of dynamic memory during interactions. Another challenge arises from the native thinking constraints of LLMs, which cannot handle complex attack combinations. We employ a question enhancement strategy based on CoT that decomposes the honeypot’s long-dialogue response tasks into three sub-tasks, allowing LLMs to assess the command’s impact on the operating system during each interaction. When posing the next question, we integrate the consolidated changes in the operating system and the user command when posing the next question to improve its understanding of complex attack combinations. The third challenge lies in that the performance of LLMs in prolonged sessions is hampered by their context length limitations and inherent forgetting mechanisms. To address this issue, we employ a Memory Pruning algorithm based on CoT that strategically prunes the session history based on its significance and temporal characteristics, ensuring that only the most essential contextual information is retained within the confines of the context length, thereby optimizing the honeypot’s effectiveness in extended interactions. The assessment of each attacker’s significance is performed by LLMs to support the automatic operation of the honeypot.
The main contributions of this work are summarized below.
@@
2. We propose a question enhancement strategy and a Memory Pruning algorithm based on the Chain of Thought (CoT) approach for terminal honeypot task scenarios. These techniques enhance the task-solving capabilities of native LLMs and balance the honeypot’s effectiveness and conversational consistency within the inherent context length limitations in extended interactions. The introduction of these strategies is essential to advancing honeypot technology and improving network security defense capabilities.
@@
4. We deploy HoneyGPT and Cowire, which is a programmatic honeypot, on the Internet for three months and observe that HoneyGPT captures more attack behaviors and longer interactions than Cowrie.
The rest of this paper is structured as follows. Section 2 provides background on LLMs, prompt engineering, and terminal honeypots. Section 3 describes the trilemma of flexibility, interaction, and deception in existing terminal honeypots. Section 4 presents our HoneyGPT solution, detailing its prompt design, strategy, workflow, and configuration. Section 5 validates the efficacy of HoneyGPT, providing baseline comparisons on deception, interaction, and flexibility, as well as field evaluations to measure its effectiveness in engaging attackers and capturing novel attack vectors. Section 6 outlines the limitations and potential future work, and finally, Section 7 concludes the paper.
II Background
@@
II-A Large Language Models
Large Language Models (LLMs) are a subset of pre-trained language models (PLMs), characterized by their voluminous parameter sets, reaching or even exceeding hundreds of billions, as defined by Kaplan et al [11]. Since OpenAI introduced ChatGPT  [19], LLMs have been widely used across a spectrum of disciplines  [6, 4], particularly in the field of cybersecurity  [16, 9, 8, 22, 21, 30]. Compared to previous models, LLMs offer two main advantages in interdisciplinary applications:
Firstly, LLMs display astonishing emergent abilities as identified by Wei et al [34], which are competencies absent in smaller-scale models but manifest in larger configurations. Such models experience a quantum leap in performance relative to their diminutive counterparts. Prominent among these emergent abilities are: In-context learning: LLMs can, upon receiving natural language instructions or several task demonstrations, generate accurate outputs for test cases by continuing the input text sequence, circumventing the need for further training or gradient adjustments  [4]. Instruction following: LLMs possess the ability to comprehend and execute natural language instructions with minimal or even zero examples, adapting to new tasks [29]. By fine-tuning tasks articulated through instructional prompts, LLMs have demonstrated proficiency in accurately handling tasks they have never seen before [20]. Step-by-step reasoning: Distinct from conventional PLMs, LLMs are capable of deconstructing complex tasks into a sequence of discrete subtasks by employing prompt-based strategies such as chain-of-thought. It is conjectured that this capacity may be derived from their training on code  [35, 6].
@@ II-B Prompt Engineering
Components. The standard components of LLM prompts typically include: 1. Task Description: Instructions that the LLM is expected to adhere to, expressed in natural language. 2. Input Data: The requisite data for a task, presented in natural language. 3. Contextual Information: The contextual or background information that can aid in the execution of a task. 4. Prompt Style: The formulation of the prompt, which is tailored to optimize the model’s responses, such as through role-playing or incremental reasoning.
@@
Design Principles. The development of effective LLM prompts is predicated on: 1. Clarity: Ensure instructions are explicit and communication is clear through symbolic delimiters, structured outputs, conditional logic, and ”Few-shot” examples. 2. Deliberation: Allow LLMs time to reason by structuring prompts for sequential information processing using ”chain of thought” techniques.
@@
Prompting Strategies. The ”Chain of Thought” (CoT)  [35, 5, 38, 13, 15, 37, 40] paradigm has been shown to significantly enhance LLM performance on tasks that require complex reasoning by incorporating intermediary reasoning steps within the prompts. The zero-shot CoT, as introduced in foundational research, employs prompts such as ”Let’s think step by step” to shepherd LLMs to a logical conclusion, thus highlighting the significance of scale in emergent capabilities.
II-C Terminal Honeypots
@@
Honeypots are security tools designed as decoys to attract and monitor unauthorized or malicious attackers. Based on their implementation approaches, terminal honeypots can be categorized into three primary types: emulated honeypots, real-system honeypots, and hybrid honeypots.
Emulated Honeypots are software constructs that simulate network services and devices to attract cyber adversaries using standard programming techniques. These honeypots range from low (Honeyd [17]) to medium (Cowrie[18]) interactivity but face constraints due to high development and operational costs, with trade-offs in adaptability and increased costs. Such frameworks have laid the groundwork for a range of specialized offshoots [14, 1].
Real-system Honeypots utilize genuine, unmodified system environments to provide a high level of interactions [33, 2, 31]. This approach enhances the detection of complex attacks but suffers from high deployment and maintenance costs and reduced adaptability due to their rigidity [1].
Hybrid Honeypots aim to merge the advantages of emulated and real-system honeypots through a scalable hybrid framework [36, 39]. This framework employs programmed methods for simulating low-level interactions, while delegating high-level interactions to real file system. While this enhances flexibility and interactivity, it also faces challenges in harmonizing simulated protocols with actual system responses, impacting their effectiveness against sophisticated attacks.
III Trilemma of Existing Terminal Honeypots
@@
The Trilemma of Terminal Honeypots reflects the critical balance required between interaction depth, flexibility, and deception capability. Honeypots from their inception have grappled with this trilemma, consistently facing trade-offs that impede their overall efficacy.
Limited Flexibility refers to the system’s adaptability in authentically replicating a variety of systems and services, which is pivotal but challenging. Emulated honeypots incur high costs and struggle with seamlessly transitioning across varied terminal environments while maintaining high interaction fidelity. Real-system honeypots lack the necessary adaptability and carry significant management overhead. The main challenge lies in designing honeypots that balance realistic interactions and adaptability without excessive costs or management burdens.
Limited Interaction is a result of limitations in implementation methods and development costs. Emulated honeypots are limited to simple rule matching and cannot execute complex command sequences like a real OS, reducing their interactive capability and effectiveness in engaging sophisticated attackers. Enhancing honeypot interaction is essential for improving their reliability and effectiveness.
Limited Deception is often due to the static and inflexible behaviors of existing honeypot systems. Emulated honeypots, restricted by static configurations and predefined rules, fail to simulate complex interactions, making them easily identifiable to experienced attackers. Similarly, real-system honeypots, despite using authentic environments, are constrained by cost and rigidity, offering limited configurations and restricted permissions that do not sufficiently engage attackers, and thus failing to entice attackers into a more profound engagement.
IV HoneyGPT
@@
To address the trilemma of existing honeypots, we present HoneyGPT, an advanced honeypot framework based on ChatGPT. HoneyGPT transforms the conventional ”request-response” message interaction (based on terminal protocols) into a ”question-answer” text interaction (via ChatGPT API).
@@
As shown in Figure 1, the framework consists of three main components: the Terminal Protocol Proxy, the Prompt Manager, and ChatGPT. The Terminal Protocol Proxy parses attackers’ commands from request messages and forwards them to the Prompt Manager. Additionally, it encapsulates terminal output received from the Prompt Manager into response messages, which are then sent back to the attackers. The Prompt Manager generates prompts to interact with ChatGPT and, based on the response, extracts the terminal output to be returned to the attackers while maintaining the state of honeypot. It employs CoT-based question enhancement and memory pruning mechanisms to improve the consistency and authenticity of the honeypot during ongoing interactions. ChatGPT generates response outputs based on the provided prompts, which are then analyzed by the Prompt Manager for further processing and state management. Since the Terminal Protocol Proxy just repurposes the work of the Cowrie honeypot, we mainly elaborate on the design of the Prompt Manager in the following.
Figure 1: HoneyGPT Framework
Figure 2: HoneyGPT workflow
IV-A The Constitution of Prompt